OpenStudy (anonymous):

Find the exact value of the trigonometric function. Do not use a calculator. \[\cot(\frac{ -5\pi }{ 4 })\] How do you do so without a calculator? Does cotangent have it's own number or something? Please explain..

OpenStudy (imstuck):

The "opposite" of cotangent is tangent. So if we are looking at the cotangent of a value, we can flip it to find the tangent. So the tangent of the above angle is \[-\frac{ 4\pi }{ 5 }\]

OpenStudy (imstuck):

Convert that to degrees so you can see what youre looking at in something other than radians. That radian measure is the same as -144 degrees.

OpenStudy (anonymous):

\[\cot \left( \frac{ -5 \pi }{ 4 } \right)=-\cot \left( \frac{ 5 \pi }{ 4 } \right)=-\cot \left( \pi+\frac{ \pi }{ 4 } \right)=-\cot \frac{ \pi }{ 4 }=-1\]

OpenStudy (anonymous):

How do you convert radians into degrees?

OpenStudy (anonymous):

\[\pi ~ radians=180 ~degree\]

OpenStudy (anonymous):

Ooooh I see. Also, where exactly did the 5 go when you did -cot(pi + (pi/4))

OpenStudy (anonymous):

take 4 L,C,M and then you will see they are equal.

OpenStudy (anonymous):

\[\pi +\frac{ \pi }{ 4 }=\frac{ 4 \pi+ \pi }{ 4 }=\frac{ 5 \pi }{ 4 }\]

OpenStudy (anonymous):

I see. Thank you @surjithayer and @IMStuck ♥

OpenStudy (anonymous):

yw